Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name brads
Log entry time 14:23:50 on March 11, 2009
Entry number 265604
Followups:
keyword=PSh/Sh logic set to OR for T6 Cosmics <-- MUST be restored before 9pm
We are re-uploading the HVs for BigBite in negative polarity and need to
take another cosmics run. The trigger has been set to Sh.OR.PSh for this
purpose.
This change must be reverted prior to beam.
A copy of this log entry has been emailed to: ellie